A bicyclist starts from rest and accelerates at a rate of 4.00 m/s2 for 3.00 s. The cyclist then travels for 35.0 s at a constant speed. How far does the cyclist travel?
Answer #1

solution Initial velocity = U= omis 4 m/s² time t acceleration = a= 3.oos SE ut + 1/ hat? Si=0+ 1x4 (3)? 4 X 9 = 18m S = 18m